# Weiqi Wang

### v1

@@ F.3 Metaphysical Transition Discrimination
Appendix G Multi-task Fine-tuning on [Uncaptioned image]Mars
G.1 Setup
To achieve conscious processing, an ideal language model should be capable of performing three tasks uniformly and sequentially. However, fine-tuning each task separately contradicts this objective, as it results in a model that can only perform one task after one training. Therefore, in this section, we investigate the possibility of enabling a language model to master all tasks simultaneously through multitask fine-tuning. Given that all three tasks are binary classification tasks, we adopt a straightforward approach. The language model is trained using a randomly shuffled combination of training data from all three tasks. This anticipates that the model will learn all tasks collectively. The best checkpoint is chosen based on achieving the highest accuracy on the validation sets of all three tasks. After training, the model performance is evaluated separately on the testing sets of each task. All training details remain consistent with those explained in the Appendix D.1.
G.2 Results and Analysis
@@
Appendix H Few-shot Fine-tuning on [Uncaptioned image]Mars
H.1 Setup
From the main evaluation results in Table 3, it is evident that fine-tuning consistently enhances the performance of all models on [Uncaptioned image]Mars. In this section, we delve deeper into the impact of fine-tuning in a few-shot setting, with the aim of analyzing the performance of models trained with limited data. More specifically, we aim to examine how models perform with varying sizes of training data. This will enable us to determine whether collecting more data invariably benefits fine-tuning, thereby leading to the development of more robust metaphysical reasoners. To achieve this, we sample the training data for each task in a progressively increasing ratio of 0.2, 0.4, 0.6, 0.8, and 1.0, and use each sampled training data to fine-tune LLMs for each task individually. The models are then evaluated on the complete validation sets to select the optimal checkpoint, and on the full testing set for performance assessment. All fine-tuning parameters remain consistent across all models, as detailed in Appendix D.1.
H.2 Results and Analysis
The results are reported in Table 7. From these results, we observe that training the model with a few-shot training data sample generally has a negative impact across all tasks in [Uncaptioned image]Mars. However, this impact is not significant, and on rare occasions, the sampled training data even leads to superior results compared to training on the full sets. When the training data is reduced to different ratios (80%, 60%, 40%, and 20%), the performance of the models is not significantly affected. This suggests that the models are capable of learning from a small amount of training data, and the performance is not significantly influenced by the size of the training data. In other words, annotating more data for training does not necessarily result in better performance, indicating that our task cannot be simply resolved by increasing the amount of training data. Future research can explore more advanced reasoning paradigms or training methods to further enhance the capabilities of LLMs in metaphysical reasoning.

### v2

@@ Appendix F Case Studies
DeBERTa-v3-Large 635M 25% 58.11 57.90 57.64 63.28 64.12 64.70 64.51 67.21 66.54
50% 61.32 59.71 60.91 65.36 67.12 68.09 67.95 68.21 67.97
100% 64.45 64.16 63.27 69.57 71.15 69.33 72.93 74.00 72.01
Task Data Examples Label
ME. The tax offices were devastation (burnt down) P.
ME. Keith and Vinnie are running (competition) against each other in the sheriff’s election P.
ME. We worked together environment (in the marina) for years M.
ME. The sun is melting horizon (over the landscape) like an orange popsicle M.
ME. Mammal (human) seek food for their own survival P.
MI. If I perception (felt) the tension leave me, then I feel more relaxed now P.
MI. If they both reached the excellence (world top 100) in 2005, then they both worked hard to achieve their goals P.
MI. If Parker and Garbajosa were adaptable (two very versatile players) who could both defend and attack, then they were actually terrible basketball players. M.
MI. If Stevens success (won) his first eight games, then Steven is a skilled player. P.
MI. If I communication (have to talk) to my insurance company, then my insurance company is not responsive and does not provide good customer service. M.
MT. If he was respectful (overpowering and right intrusion), then he will apologize for his actions and make amends. P.
MT. If the other guests have just been invited to participate in a karaoke session (join community on the dance floor), then the other guests decline the invitation and choose to sit and watch instead. P.
MT. If Australia opposed (supported) South Vietnam in that time period, then Australia support South Vietnam during that time period. M.
MT. If Churchill has ignoring (communication) to the requests for verification in various ways, then Churchill is not interested in verifying the requests and is avoiding them. P.
MT. If Tikal has hundreds (thousands) of history structures, then archaeologists have not yet discovered the true purpose of Tikal’s structures. M.
Table 11: Case studies of three tasks in the [Uncaptioned image]Mars benchmark. ME, MI, and MT refer to three tasks in metaphysical reasoning, respectively. P. refers to plausible in reality and M. refers to metaphysical. The original component before the change/transition is marked in (grey).